14. Comfort realty has 100,000 shares outstanding with a market value of $150/share and a par value of $10 per share. What is the number of shares outstanding and the value of each share when the company does a 3-for-1 stock split?
Answer #1

Number of shares outstanding after stock split= 300000

Market value per share= $50

Working 1

Before split

After split

Number of shares outstanding

100000

300000

Market value per share

$                  150.00

$                      50.00

Total market value of shares or Market cap

$ 1,50,00,000.00

$    1,50,00,000.00

When stock split is done the total stockholder’s equity remains same only the number of shares increase.

Working 2

Before split

After split

Number of shares outstanding

100000

=100000*3

Market value per share

150

=150/3

Total market value of shares or Market cap

=100000*150

=300000*50